# Distance Between Stochastic Linear Systems

### v1

@@ Abstract
This manuscript proposes a distance measure between stochastic linear dynamical systems. While the existing stochastic control theory is well equipped to handle dynamical systems with stochastic uncertainties, a paradigm shift using distance measure based decision making is required for the effective further exploration of the field. As a first step, a distance measure between two linear time invariant stochastic dynamical systems is proposed here, extending the existing distance metrics between deterministic linear dynamical systems. Distance measure for stochastic systems is proposed for the frequency domain setting as the worst-case point-wise in frequency Wasserstein distance between distributions characterising the uncertainties using inverse stereographic projection on the Riemann sphere. For the time domain setting, the proposed distance corresponds to the gap metric induced type-$q$ Wasserstein distance between the push-forward measures under both systems’ corresponding measurable maps from the parameter space to their respective space of system plants. It is proved and demonstrated using numerical simulation that the proposed frequency domain distance measure shall never exceed the proposed time domain distance measure counterpart. Lower and upper bounds are provided for the proposed distance measures in both frequency and time domain settings. The proposed distance measures induce a topology in the corresponding (frequency/time) domain space of stochastic dynamical systems and will facilitate the provision of probabilistic guarantees on system robustness and controller performances.
keywords:
@@ 1 Introduction
Contributions: The main contributions of this paper are as follows:
@@
    In the frequency domain setting, the proposed distance measure given by (3.6) refers to the worst-case point-wise-in-frequency type-$q$ chordal metric induced Wasserstein distance between distributions governing the uncertainties of the two stochastic linear systems in the Riemann sphere. Under the assumption of distributions being uniform in nature, a support distance based upper bound for proposed distance measure is given in Theorem˜3.9. On the other hand, using the deviation of the perturbed models of each systems from their respective nominal models, a lower bound for the proposed distance measure is given in Theorem˜3.11. Extensions with respect to empirical distribution case are also given in Proposition˜3.14 and Theorem˜3.18.
@@
    In the time domain setting, the proposed distance measure given by (48) refers to the gap metric induced type-$q$ Wasserstein distances between the distributions governing the uncertainty of the systems obtained through the push-forward of the distribution of uncertain system parameters under the measurable mapping that connects the parameters and the perturbed model. Upper bounds for the proposed distance measure are proposed in Proposition˜4.4 and Theorem˜4.6. On the similar lines of the frequency domain setting, a lower bound for the proposed distance measure in the time domain is given using the deviation of the perturbed models of each systems from their respective nominal models in Proposition˜4.8.
@@
    We also prove in Theorem˜5.5 that for stochastic LTI systems, the proposed frequency domain distance never exceeds the time domain distance, mimicking the inequality relationship that exists between the $\nu$-gap metric and the gap metric in the deterministic systems setting.
@@
Following a detailed discussion on the notations and the preliminaries needed for problem formulation in both frequency and time domain settings in Section˜2, we will first begin our problem formulation in the frequency domain setting and propose a valid distance metric in Section˜3. Following that, we will present the analogous problem formulation in the time domain setting and propose an analogous valid distance metric in Section˜4. The proposed distance measures in the frequency and time domain settings are compared in Section˜5. Finally, the paper is closed in Section˜6 along with the summary of findings and directions for future research. Simulation results are provided throughout the paper to demonstrate the proposed concepts and guarantees. All the Matlab codes responsible for reproducing the simulation results provided in the paper can be found at https://github.com/venkatramanrenganathan/stochastic-system-gap.
@@ 2.1 Function Spaces & Norms
(1) $\left\|P\right\|_{\mathcal{H}_{\infty}}:=\sup_{\begin{subarray}{c}u\in\mathcal{H}_{2}\\ u\neq 0\end{subarray}}\frac{\left\|Pu\right\|_{\mathcal{H}_{2}}}{\left\|u\right\|_{\mathcal{H}_{2}}}.$
@@
(2) $\left\|P\right\|_{\mathcal{L}_{\infty}}:=\mathop{\mathrm{ess\,sup}}_{\omega\in\mathbb{R}}\bar{\sigma}(P(j\omega)).$
@@
(3) $W^{q}_{q}(\mathbb{Q}_{1},\mathbb{Q}_{2})\overset{\Delta}{=}\underset{\pi\in\Pi(\mathbb{Q}_{1},\mathbb{Q}_{2})}{\inf}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left\|z_{1}-z_{2}\right\|^{q}\pi(dz_{1},dz_{2}).$
@@
(4) $(f_{\#}\mu)(\mathbf{B}):=\mu\left(f^{-1}(\mathbf{B})\right),$
@@
(5) $f^{-1}(\mathbf{B}):=\left\{x\in\mathcal{X}\mid f(x)\in\mathcal{B}\right\}.$
@@ 2.2 Preliminaries on Stereographic Projections
(6) $\mathfrak{R}=\left\{(x,y,z)\in\mathbb{R}^{3}\mid x^{2}+y^{2}+\left(z-0.5\right)^{2}=\left(0.5\right)^{2}\right\}.$
@@
(7) $\mathfrak{R}=\left\{(r,\theta,\varphi)\in\mathbb{R}^{3}\mid r=\frac{1}{2},\ \theta\in[0,\pi],\ \varphi\in[0,2\pi]\right\}.$
@@
(8) $c=\phi(R):=\left(\frac{r_{x}}{1-r_{z}}\right)+j\left(\frac{r_{y}}{1-r_{z}}\right).$
@@
(9) $(r_{x},r_{y},r_{z})=\left(\frac{\mathrm{Re}(c)}{1+\mid{c}\mid^{2}},\frac{\mathrm{Im}(c)}{1+\mid{c}\mid^{2}},\frac{\mid{c}\mid^{2}}{1+\mid{c}\mid^{2}}\right).$
@@
(10) $\mathbb{P}_{R}$ $=\frac{r+r^{3}}{2}\mathbb{P}_{\mathsf{xy}}(x,y),\quad\text{where }r=\sqrt{x^{2}+y^{2}}.$
@@ 2.3 Preliminaries on Gap Metric
(12a) $\mathcal{G}_{P}$ $:=\left\{(u,y)\mid y=Pu\right\}\subseteq\mathcal{H}_{2}\times\mathcal{H}_{2}$
(12b) $=\underbrace{\begin{bmatrix}D\\ N\end{bmatrix}}_{=:G}\mathcal{H}_{2}=\mathrm{Range}(G),$
@@
(13) $\delta_{g}(P_{1},P_{2})=\left\|\Pi_{\mathcal{G}_{1}}-\Pi_{\mathcal{G}_{2}}\right\|.$
@@
(14) $b_{P,C}=\left\|\begin{bmatrix}P\\ I\end{bmatrix}(I-CP)^{-1}\begin{bmatrix}-C&I\end{bmatrix}\right\|^{-1}_{\mathcal{H}_{\infty}}$
@@ 3 Problem Formulation in Frequency Domain
(15) $P_{i}(\theta_{i};s)=C_{i}(\theta_{i})(sI-A_{i}(\theta_{i}))^{-1}B_{i}(\theta_{i}),$
@@
(16) $\mathbb{P}_{P_{i}(\omega)}:=\left(\Phi^{\mathbb{C}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}},\quad\text{where}\quad\Phi^{\mathbb{C}}_{i}=\mathcal{E}_{\omega}\circ P_{i},$
@@
(17) $\left(\Phi^{\mathbb{C}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}(\mathbf{B})=\mathbb{P}(\Phi^{\mathbb{C}}_{i}\in\mathbf{B})=\mathbf{f}_{\theta_{i}}\left(\left(\Phi^{\mathbb{C}}_{i}\right)^{-1}(\mathbf{B})\right).$
Assumption 2.
@@
(18) $\mathcal{R}_{P_{\ell}}$ $=\phi^{-1}(\mathcal{S}_{P_{\ell}}):=\left\{R\in\partial\mathfrak{R}\mid R=\phi^{-1}(r),\forall r\in\mathcal{S}_{P_{\ell}}\right\}.$
@@ 3.1 Stereographic Projection of Distribution
(19) $\mathbb{P}_{R_{i}(\omega)}:=\left(\Phi^{\mathfrak{R}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}},\quad\text{where}\quad\Phi^{\mathfrak{R}}_{i}=\phi^{-1}\circ\Phi^{\mathbb{C}}_{i},$
@@
(20) $\left(\Phi^{\mathfrak{R}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}(\mathbf{B})=\mathbb{P}(\Phi^{\mathfrak{R}}_{i}\in\mathbf{B})=\mathbf{f}_{\theta_{i}}\left(\left(\Phi^{\mathfrak{R}}_{i}\right)^{-1}(\mathbf{B})\right).$
@@ 3.2 Support Distance Between Systems $P_{1}$ and $P_{2}$
(21) $\mathrm{d^{\mathbb{C}}_{sup}}(P_{1},P_{2})$ $:=\sup_{\omega\in\Omega}\,\sup_{\begin{subarray}{c}s_{1}\in\mathcal{S}_{P_{1}}(\omega)\\ s_{2}\in\mathcal{S}_{P_{2}}(\omega)\end{subarray}}\left\|s_{1}-s_{2}\right\|.$
@@
Given two points $r_{1}\in\mathcal{R}_{P_{1}}(\omega)$ and $r_{2}\in\mathcal{R}_{P_{2}}(\omega)$ and $\mathfrak{R}_{c}=(0,0,0.5)$, the geodesic distance between them, denoted by $\mathrm{d_{geo}}(r_{1},r_{2})$ and the chordal distance between them, denoted by $\mathrm{d_{chord}}(r_{1},r_{2})$ are respectively given by
(22) $\mathrm{d_{geo}}(r_{1},r_{2})$ $:=\frac{1}{2}\cos^{-1}\left(4\langle(r_{1}-\mathfrak{R}_{c}),(r_{2}-\mathfrak{R}_{c})\rangle\right),\quad\text{and}$
(23) $\mathrm{d_{chord}}(r_{1},r_{2})$ $:=\left\|r_{1}-r_{2}\right\|=\sin(\mathrm{d_{geo}}(r_{1},r_{2})),\quad\text{and}$
(24) $\mathrm{d_{chord}}(r_{1},r_{2})$ $\leq\mathrm{d_{geo}}(r_{1},r_{2}).$
Proof 3.4.
For the Riemann sphere centred at $\mathfrak{R}_{c}=(0,0,0.5)$ and of radius $R=0.5$, let $\theta$ be the central angle measured at $\mathfrak{R}_{c}$ between the points $r_{1}\in\mathcal{R}_{P_{1}}(\omega)$ and $r_{2}\in\mathcal{R}_{P_{2}}(\omega)$. Then, $\mathrm{d_{geo}}(r_{1},r_{2})=R\theta$ which evaluates to (22). Note that $\mathrm{d_{geo}}(r_{1},r_{2})=R\theta\iff\theta=\frac{\mathrm{d_{geo}}(r_{1},r_{2})}{R}=2\mathrm{d_{geo}}(r_{1},r_{2})$. Similarly, $\mathrm{d_{chord}}(r_{1},r_{2})=2R\sin\left(\frac{\theta}{2}\right)$ which then evaluates to (23). Since $\sin(x)\leq x,\forall x\geq 0$, the result (24) follows immediately.
@@
Given two systems $P_{\ell}$ with $\ell=\{1,2\}$, whose uncertainty in $\mathbb{C}$ are characterized by distributions defined on support sets $\mathcal{S}_{P_{\ell}}(\omega)$ for every frequency $\omega\in\Omega$, we define the support distance in the Riemann sphere between the systems as the worst-case pointwise-in-frequency chordal distance between points in their projected support sets $\mathcal{R}_{P_{\ell}}(\omega)$. That is,
(25) $\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})$ $:=\sup_{\omega\in\Omega}\,\underbrace{\sup_{\begin{subarray}{c}r_{1}\in\mathcal{R}_{P_{1}}(\omega)\\ r_{2}\in\mathcal{R}_{P_{2}}(\omega)\end{subarray}}\mathrm{d_{chord}}(r_{1},r_{2})}_{:=\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2},\omega)}.$
Clearly, both the support distances $\mathrm{d^{\mathbb{C}}_{sup}}(P_{1},P_{2})$ and $\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})$ given by (21) and (25) do not explicitly take into account the information of the distributions $\mathbb{P}_{P_{1}(\omega)}$ and $\mathbb{P}_{P_{2}(\omega)}$ and rather is based only on their support sets $\mathcal{S}_{P_{1}}(\omega)$ and $\mathcal{S}_{P_{2}}(\omega)$ respectively for every frequency $\omega\in\Omega$. That is, both the support distances $\mathrm{d^{\mathbb{C}}_{sup}}(P_{1},P_{2})$ and $\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})$ given by (21) and (25) do not take into account the frequency with which each plant models occur within their respective support sets meaning that it just encodes the physical separation between the plant models getting realised in the respective support sets. This shortcoming can be addressed by defining a chordal metric induced Wasserstein distribution between distributions living in the Riemann sphere which will be shortly.

### v2

@@ Abstract
While the existing stochastic control theory is well equipped to handle dynamical systems with stochastic uncertainties, a paradigm shift using distance measure based decision making is required for the effective further exploration of the field. As a first step, a distance measure between two stochastic linear time invariant systems is proposed here, extending the existing distance metrics between deterministic linear dynamical systems. In the frequency domain, the proposed distance measure corresponds to the worst-case point-wise in frequency Wasserstein distance between distributions characterising the uncertainties using inverse stereographic projection on the Riemann sphere. For the time domain setting, the proposed distance corresponds to the gap metric induced type-q Wasserstein distance between the distribution characterising the uncertainty of plant models. Apart from providing lower and upper bounds for the proposed distance measures in both frequency and time domain settings, it is proved that the former never exceeds the latter. The proposed distance measures will facilitate the provision of probabilistic guarantees on system robustness and controller performances.
{IEEEkeywords}
@@ 1 Introduction
Contributions: The main contributions of this paper are:
@@
    In the frequency domain setting, the proposed distance measure given by (3.3) refers to the worst-case point-wise-in-frequency type-$q$ chordal metric induced Wasserstein distance between distributions governing the uncertainties of the two stochastic linear systems in the Riemann sphere. Under the assumption of distributions being uniform in nature, a support distance based upper bound for proposed distance measure is given in Theorem 3.2. On the other hand, using the deviation of the perturbed models of each systems from their respective nominal models, a lower bound for the proposed distance measure is given in Theorem 3.3. Extensions with respect to empirical distribution case are also given in Proposition 3.4 and Theorem 3.6.
@@
    In the time domain setting, the proposed distance measure given by (49) refers to the gap metric induced type-$q$ Wasserstein distances between the distributions governing the uncertainty of the systems obtained through the push-forward of the distribution of uncertain system parameters under the measurable mapping that connects the parameters and the perturbed model. Upper bounds for the proposed distance measure are proposed in Proposition 4.2 and Theorem 4.3. On the similar lines of the frequency domain setting, a lower bound for the proposed distance measure in the time domain is given using the deviation of the perturbed models of each systems from their respective nominal models in Proposition 4.4.
@@
    We also prove in Theorem 5.3 that for stochastic LTI systems, the proposed frequency domain distance never exceeds the time domain distance, mimicking the inequality relationship that exists between the $\nu$-gap metric and the gap metric in the deterministic systems setting.
@@
Following a detailed discussion on the notations and the preliminaries needed for problem formulation in both frequency and time domain settings in Section 2, we will first begin our problem formulation in the frequency domain setting and propose a valid distance metric in Section 3. Following that, we will present the analogous problem formulation in the time domain setting and propose an analogous valid distance metric in Section 4. The proposed distance measures in the frequency and time domain settings are compared in Section 5. Finally, the paper is closed in Section 6 along with the summary of findings and directions for future research. Simulation results are provided throughout the paper to demonstrate the proposed concepts and guarantees. All the Matlab codes responsible for reproducing the simulation results provided in the paper can be found at https://github.com/venkatramanrenganathan/stochastic-system-gap.
@@ 2.1 Function Spaces & Norms
$\left\|P\right\|_{\mathcal{H}_{\infty}}:=\sup_{\begin{subarray}{c}u\in\mathcal{H}_{2}\\ u\neq 0\end{subarray}}\frac{\left\|Pu\right\|_{\mathcal{H}_{2}}}{\left\|u\right\|_{\mathcal{H}_{2}}}.$ (1)
@@
$\left\|P\right\|_{\mathcal{L}_{\infty}}:=\mathop{\mathrm{ess\,sup}}_{\omega\in\mathbb{R}}\bar{\sigma}(P(j\omega)).$ (2)
@@
$W^{q}_{q}(\mathbb{Q}_{1},\mathbb{Q}_{2})\overset{\Delta}{=}\underset{\pi\in\Pi(\mathbb{Q}_{1},\mathbb{Q}_{2})}{\inf}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\left\|z_{1}-z_{2}\right\|^{q}\pi(dz_{1},dz_{2}).$ (3)
@@
$(f_{\#}\mu)(\mathbf{B}):=\mu\left(f^{-1}(\mathbf{B})\right),$ (4)
@@
$f^{-1}(\mathbf{B}):=\left\{x\in\mathcal{X}\mid f(x)\in\mathcal{B}\right\}.$ (5)
@@ 2.2 Preliminaries on Stereographic Projections
$\mathfrak{R}=\left\{(x,y,z)\in\mathbb{R}^{3}\mid x^{2}+y^{2}+\left(z-0.5\right)^{2}=\left(0.5\right)^{2}\right\}.$ (6)
@@
$\mathfrak{R}=\left\{(r,\theta,\varphi)\in\mathbb{R}^{3}\mid r=\frac{1}{2},\ \theta\in[0,\pi],\ \varphi\in[0,2\pi]\right\}.$ (7)
@@
$c=\phi(R):=\left(\frac{r_{x}}{1-r_{z}}\right)+j\left(\frac{r_{y}}{1-r_{z}}\right).$ (8)
@@
$(r_{x},r_{y},r_{z})=\left(\frac{\mathrm{Re}(c)}{1+\mid{c}\mid^{2}},\frac{\mathrm{Im}(c)}{1+\mid{c}\mid^{2}},\frac{\mid{c}\mid^{2}}{1+\mid{c}\mid^{2}}\right).$ (9)
@@
$\mathbb{P}_{R}$ $=\frac{r+r^{3}}{2}\mathbb{P}_{\mathsf{xy}}(x,y),\quad\text{where }r=\sqrt{x^{2}+y^{2}}.$ (10)
@@ 2.3 Preliminaries on Gap Metric
$\mathcal{G}_{P}$ $:=\left\{(u,y)\mid y=Pu\right\}\subseteq\mathcal{H}_{2}\times\mathcal{H}_{2}$ (12a)
$=\underbrace{\begin{bmatrix}D\\ N\end{bmatrix}}_{=:G}\mathcal{H}_{2}=\mathrm{Range}(G),$ (12b)
@@
$\delta_{g}(P_{1},P_{2})=\left\|\Pi_{\mathcal{G}_{1}}-\Pi_{\mathcal{G}_{2}}\right\|.$ (13)
@@
$b_{P,C}=\left\|\begin{bmatrix}P\\ I\end{bmatrix}(I-CP)^{-1}\begin{bmatrix}-C&I\end{bmatrix}\right\|^{-1}_{\mathcal{H}_{\infty}}$ (14)
@@ 3 Problem Formulation in Frequency Domain
$P_{i}(\theta_{i};s)=C_{i}(\theta_{i})(sI-A_{i}(\theta_{i}))^{-1}B_{i}(\theta_{i}),$ (15)
@@
$\mathbb{P}_{P_{i}(\omega)}:=\left(\Phi^{\mathbb{C}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}},\quad\text{where}\quad\Phi^{\mathbb{C}}_{i}=\mathcal{E}_{\omega}\circ P_{i},$ (16)
@@
$\left(\Phi^{\mathbb{C}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}(\mathbf{B})=\mathbb{P}(\Phi^{\mathbb{C}}_{i}\in\mathbf{B})=\mathbf{f}_{\theta_{i}}\left(\left(\Phi^{\mathbb{C}}_{i}\right)^{-1}(\mathbf{B})\right).$ (17)
Assumption 3.2.
@@
$\mathcal{R}_{P_{\ell}}\!=\!\phi^{-1}(\mathcal{S}_{P_{\ell}})\!=\!\left\{\!R\in\partial\mathfrak{R}\mid R=\phi^{-1}(r),\forall r\in\mathcal{S}_{P_{\ell}}\!\right\}.$ (18)
@@ 3.1 Stereographic Projection of Distribution
$\mathbb{P}_{R_{i}(\omega)}:=\left(\Phi^{\mathfrak{R}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}},\quad\text{where}\quad\Phi^{\mathfrak{R}}_{i}=\phi^{-1}\circ\Phi^{\mathbb{C}}_{i},$ (19)
@@
$\left(\Phi^{\mathfrak{R}}_{i}\right)_{\#}\mathbf{f}_{\theta_{i}}(\mathbf{B})=\mathbb{P}(\Phi^{\mathfrak{R}}_{i}\in\mathbf{B})=\mathbf{f}_{\theta_{i}}\left(\left(\Phi^{\mathfrak{R}}_{i}\right)^{-1}(\mathbf{B})\right).$ (20)
@@ 3.2 Support Distance Between Systems $P_{1}$ and $P_{2}$
$\mathrm{d^{\mathbb{C}}_{sup}}(P_{1},P_{2})$ $:=\sup_{\omega\in\Omega}\,\sup_{\begin{subarray}{c}s_{1}\in\mathcal{S}_{P_{1}}(\omega)\\ s_{2}\in\mathcal{S}_{P_{2}}(\omega)\end{subarray}}\left\|s_{1}-s_{2}\right\|.$ (21)
@@
Given two points $r_{1}\in\mathcal{R}_{P_{1}}(\omega)$, and $r_{2}\in\mathcal{R}_{P_{2}}(\omega)$, the geodesic distance between them, denoted by $\mathrm{d_{geo}}(r_{1},r_{2})$ and the chordal distance between them, denoted by $\mathrm{d_{chord}}(r_{1},r_{2})$ are respectively given by
$\mathrm{d_{geo}}(r_{1},r_{2})$ $:=\frac{1}{2}\cos^{-1}\left(4\langle(r_{1}-\mathfrak{R}_{c}),(r_{2}-\mathfrak{R}_{c})\rangle\right),$ (22)
$\mathrm{d_{chord}}(r_{1},r_{2})$ $:=\left\|r_{1}-r_{2}\right\|=\sin(\mathrm{d_{geo}}(r_{1},r_{2})),\quad\text{and}$ (23)
$\mathrm{d_{chord}}(r_{1},r_{2})$ $\leq\mathrm{d_{geo}}(r_{1},r_{2}).$ (24)
Proof.
For the Riemann sphere centred at $\mathfrak{R}_{c}=(0,0,0.5)$ and of radius $R=0.5$, let $\theta$ be the central angle measured at $\mathfrak{R}_{c}$ between the points $r_{1}\in\mathcal{R}_{P_{1}}(\omega)$ and $r_{2}\in\mathcal{R}_{P_{2}}(\omega)$. Then, $\mathrm{d_{geo}}(r_{1},r_{2})=R\theta$ which evaluates to (22). Note that $\mathrm{d_{geo}}(r_{1},r_{2})=R\theta\iff\theta=\frac{\mathrm{d_{geo}}(r_{1},r_{2})}{R}=2\mathrm{d_{geo}}(r_{1},r_{2})$. Similarly, $\mathrm{d_{chord}}(r_{1},r_{2})=2R\sin\left(\frac{\theta}{2}\right)$ which then evaluates to (23). Since $\sin(x)\leq x,\forall x\geq 0$, the result (24) follows. ∎
@@
Given two systems $P_{\ell}$ with $\ell=\{1,2\}$, whose uncertainty in $\mathbb{C}$ are characterized by distributions defined on support sets $\mathcal{S}_{P_{\ell}}(\omega)$ for every frequency $\omega\in\Omega$, the support distance in the Riemann sphere between the systems is defined as the worst-case pointwise-in-frequency chordal distance between points in their projected support sets $\mathcal{R}_{P_{\ell}}(\omega)$. That is,
$\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})$ $:=\sup_{\omega\in\Omega}\,\underbrace{\sup_{\begin{subarray}{c}r_{1}\in\mathcal{R}_{P_{1}}(\omega)\\ r_{2}\in\mathcal{R}_{P_{2}}(\omega)\end{subarray}}\mathrm{d_{chord}}(r_{1},r_{2})}_{:=\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2},\omega)}.$ (25)
Clearly, both the support distances $\mathrm{d^{\mathbb{C}}_{sup}}(P_{1},P_{2})$ and $\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})$ given by (21) and (25) do not explicitly take into account the information of the distributions $\mathbb{P}_{P_{1}(\omega)}$ and $\mathbb{P}_{P_{2}(\omega)}$ and rather is based only on their support sets $\mathcal{S}_{P_{1}}(\omega)$ and $\mathcal{S}_{P_{2}}(\omega)$ respectively for every frequency $\omega\in\Omega$. That is, both the support distances $\mathrm{d^{\mathbb{C}}_{sup}}(P_{1},P_{2})$ and $\mathrm{d^{\mathfrak{R}}_{sup}}(P_{1},P_{2})$ given by (21) and (25) do not take into account the frequency with which each plant models occur within their respective support sets meaning that it just encodes the physical separation between the plant models getting realised in the respective support sets. This shortcoming can be addressed by defining a chordal metric induced Wasserstein distribution between distributions living in the $\partial\mathfrak{R}$.